# Tailoring the Hyperparameters of a Wide-Kernel Convolutional Neural Network to Fit Different Bearing Fault Vibration Datasets

### v1

@@ 3.4 Summary of Datasets Used
Across the various parts of our research, multiple datasets are employed. For the sake of clarity we now present the full collection of datasets we use and remark on which architectures they are used with. Table 4 shows these details.
@@
Table 4: Summary of datasets used in the research.
4 Results
@@
4.1 Results Pt. 1 – Experiments on Multiple Architectures
The overall focus of our work is a specific wide-kernel CNN architecture. However, it is important to note that other architectures also present the user with hyperparameter decisions, and also suffer from the fact that poor hyperparameter choices can be disastrous for accuracy. We therefore briefly present some experiments with two other popular architectural styles: a recurrent LSTM type of neural network and a transformer network. These results establish the general importance of hyperparameter tuning when working with neural networks, which is the basis for the remaining sections where we deal with the wide-kernel architecture much more extensively.
4.1.1 Long Short Term Memory (LSTM)
Across our sample of 100 LSTM configurations, we see that poor hyperparameters can be disastrous. In the worst cases (CWRU and XJTU) the lowest-performing network was wrong more than 10 times as often as the best network. On all datasets the default values do roughly as well as about 75% of modified hyperparameter settings and therefore are reasonable choices. One interesting thing is that the LSTM network’s default values appear to under-perform on XJTU data. Facing this benchmark, an LSTM configured with the defaults only achieved an accuracy of 72% whilst it was possible to get above 90% with other hyperparameter values. Re-tuning the default hyperparameter settings seems to be a good idea on XJTU data, indicating that hyperparameters that perform well on one dataset do not necessarily perform well on another dataset. The default values fared relatively much better on CWRU and SEU. Overall the results show that hyperparameter tuning has a noticeable impact on the LSTM architecture much like with the transformer and wide-kernel CNN networks.
Min 25% Median 75% Max w/ Defaults
@@ 4.1.1 Long Short Term Memory (LSTM)
Table 5: Tukey’s five-number summary of how accurate the LSTM architecture was with different hyperparameter settings. The accuracy score when using the default hyperparameter values is also shown in the right-hand column.
4.1.2 Transformer
The transformer architecture performs relatively well when using the default hyperparameter settings, as shown in Table 6. On four of the baselines it was able to correctly identify fault conditions on more than 90% of test cases. Looking across all of the baseline datasets, the default hyperparameter settings were at least as good as 75% of altered versions of the settings. The results suggest that the default values are quite reasonable choices. On the Gearbox dataset the transformer was inaccurate no matter if the default hyperparameter settings were used or not, suggesting a specific problem with that one dataset. Importantly, our results show overall that the transformer architecture is sensitive to hyperparameter tweaking. On CWRU and UOC for example, the accuracy plunges by roughly 90% when swapping the best-performing hyperparameters for the worst.
@@ 4.1.2 Transformer
Table 6: Tukey’s five-number summary of how accurate the transformer architecture was with different hyperparameter settings. The accuracy score when using the default hyperparameter values is also shown in the right-hand column.
4.2 Results Pt. 2 – Focusing on the Wide-Kernel CNN Architecture
The main body of our research is about the wide-kernel CNN architecture described in Section 3.3.1. We begin the results by revisiting how successful networks in a grid search are when trained and tested on seven different benchmark datasets. This grid search is the same as was analysed before in [1], although here our results have a different focus. We look at the average test results when setting individual hyperparameters to specific values, and we also zoom in on the interactions between the hyperparameters that were previously identified as most important. These results help us formulate guidance on how to tune the hyperparameters of the wide-kernel CNN, in later sections of this paper.
@@
4.2.1 Summarising Grid Search Results
Now we look at how the accuracy of the wide-kernel CNN varies as we change both the hyperparameter values and the data. Each condition’s accuracy is shown in Table 7. These results are based on the same grid search used previously in [1], which trained and tested different combinations of hyperparameters on 7 benchmarks. We revisit the grid search in order to pull out and highlight helpful information when thinking about how to tune the hyperparameters on different datasets.
@@ 4.2.1 Summarising Grid Search Results
Kernel size layer 1 16 25 78 47 74 87 32 64
32 31 78 48 84 80 35 59
@@
Kernel size layer 2 3 43 78 50 85 71 43 67
6 45 77 49 84 70 43 62
Filters layer 2 8 44 78 52 85 74 49 68
16 50 78 52 85 73 50 66
@@
Kernel size layers 3-5 3 45 80 50 84 71 43 65
6 43 75 49 85 69 42 64
Filters layers 3-5 8 33 83 49 86 67 49 75
16 45 91 50 89 72 51 82
@@
Table 7: The average performance for different values of wide-kernel CNN hyperparameters when trained and tested on different datasets.
@@
Perhaps more interesting are the individual accuracy scores showing what happens when we tune different hyperparameters. The kernel size in the first layer reveals the most unexpected results. With CWRU, Paderborn and UOC there is a strong effect whereby a longer kernel does better and the longest kernel exceeds the shortest by at least 15% accuracy. The MFPT and XJTU benchmarks exhibit a weaker effect of roughly 6% difference between the longest and shortest kernel. The kernel size in layer 1 barely effects the results on Gearbox data. Meanwhile, a strong contradictory effect is observed when using SEU data. The shortest kernel is much better (more than 30% better) compared to the longest kernel size. In all, it seems that longer kernels in the first layer tend to be best but it very much depends on the fault vibration data.
@@
The number of filters in the first layer was also identified as important by past work. Good performance is often achieved when there are 16 or 32 filters in the first layer. Either 16 or 32 was the best on every dataset except XJTU (when 8 filters was 1% better) and Gearbox. With Gearbox, the trend is that more filters gives higher accuracy and 256 beats 8 filters by about 7% accuracy. Looking across all the datasets, 16 or 32 are the safest bets when tuning this hyperparameter but exceptions are possible.
The final ‘important’ hyperparameter is the number of filters in layer 3 to 5. We observe that either 16 or 32 works best on all the benchmark datasets. This conclusion at least is straightforward.
4.2.2 Influence of Important Hyperparameters on One Another
@@
We now take another deep dive into the three hyperparameters identified as important by our previous work. It is very computationally expensive to run a full grid search over all the possible hyperparameter combinations and so someone applying the network to new data might well want to know how to tune the hyperparameters one-at-a-time. We ask which is the best order to tune the three most important hyperparameters sequentially. The method we used in our previous paper [1] comes in handy again here: we look at how much one hyperparameter influences another hyperparameter which has already been tuned. Because the hyperparameters are interconnected it is possible that a hyperparameter that started off being optimal no longer has the best value when a different hyperparameter changes. We look at how often this phenomenon happens and we summarise the information as an ‘influence’ score. If A is highly influential on B then changing A most likely means changing the best value for B. We can see the numbers in Figure 6.
@@
The kernel size in layer 1 appears to be the most influential and therefore it appears to be the best one to tune first. Tuning it later on would likely disrupt the other hyperparameters and cause them to need re-tuning. In other words, there is often no point in tuning the number of filters in layers 1 or 3-5 beforehand, since they will need to be tuned again when the kernel size changes. After choosing the size of filters in the first layer it is best to choose how many next. The choice of how many filters in layers 3 to 5 can be delayed to the end.
4.2.3 Resampling
@@
The results of a grid search over the three most important hyperparameters, performed once each for the different resampling conditions, are brought together in the results here. Visualising the results makes it possible to spot any obvious trends in how the (increasing levels of) resampling impact the performance of different hyperparameter settings. We present several analyses on both CWRU 48 kHz data to see how much each hyperparameter seems to influence performance. These analyses include using a method based on Shapley values and using a method based on d-dimensional earth mover’s scores to pinpoint important hyperparameters. Here, we include the results obtained when resampling CWRU data, starting from data originally sampled at 48 kHz.
@@
Figure 7: Correlation between different resampling conditions, starting from 48 kHz.
@@
First, we consider how much the accuracy scores change as a consequence of resampling the data. Correlations tell us whether the best-performing combinations of hyperparameters on one version of the data also remain the best after resampling, and vice versa with the worst performers. The precise correlation results are shown in Figure 7. All correlations are positive which implies that the hyperparameter values that do well on one version of the data have a tendency to perform well on data resampled to a different rate. The weakest correlation is between the original 48 kHz data and the most resampled data of 3 kHz. This is not surprising since they are the furthest away from each other in terms of data properties. The weakened correlation between 48 kHz and 3 kHz gives us an initial
@@
Before moving on we must make a brief disclaimer about the correlations, in particular we acknowledge there is a ceiling effect because many networks achieve close to the maximum accuracy of 1.0, and, the training and testing of neural networks is stochastic and so there may be some random noise in the accuracy scores. There is a chance that these statistical quirks have inflated or deflated the correlations, and so we do not report p-values which might be biased by these effects. Instead, we view the results as indicative. The results give an indication that hyperparameter settings that are stronger on one version of the data are more likely to be stronger when also used on a resampled version of the data, but room is also left for differences between sampling rates to emerge.
@@
Figure 9: Feature importance after resampling to different rates, starting from 48 kHz. Each row is normalised to show the relative importance of the hyperparameters.
In the next piece of our analysis we look at feature importance. This shows us the impact of resampling on the other two most important hyperparameters alongside kernel size in the first layer. We interpret importance by saying that if a hyperparameter is important then it determines whether a neural network is accurate or not. If a hyperparameter is important and you know its value, then you can predict the accuracy of the network. We arrive at numeric importance scores by looking at the contribution (Shapley value) of each hyperparameter to the successful predictions of accuracy of a multilayer perceptron. Figure 9 shows the Shapley values after normalising each row to add up to 1.0. Naturally, higher values indicate more importance.
No matter the degree of resampling, the number of filters in layers 3-5 is the most important. Interestingly, this hyperparameter is increasingly dominant as the resampling gets more extreme. With no resampling or with resampling to 24 kHz, the other hyperparameters constitute almost half of the Shapley values. There is no obvious pattern to the kernel size, and the much higher importance when resampling to 24 kHz is an unexplained outlier. In any case, the kernel size did not keep on becoming more important as a result of escalating resampling. Instead, the clearest conclusion is that the filters in the later layers become more important when resampling is greater. If we consider that the first layer is often used to process raw, noisy signals then the declining importance of this layer might reflect the fact that more resampled data is already more processed before it reaches the neural network as an input. Alternatively, networks’ overall lower accuracies on highly resampled data (shown in Figure 8) present the possibility that resampling deletes patterns in the signal, hence the first layer is important when processing mildly resampled data since it allows networks to pick up on patterns useful for successfully classifying vibrations.
Figure 10: How much the accuracy scores change as a result of tweaking a hyperparameter, when processing data with different levels of resampling, starting from 48 kHz. Each cell is normalised to show its value relative to surrogate data (see explanation in Section 3.3.6).
Another way to consider how important different hyperparameters are is to ask how much the accuracy scores change when the hyperparameter is changed. The d-dimensional earth mover’s distance (EMD) quantifies just this. Figure 10 shows how many times larger the observed EMD is than a surrogate baseline. Firstly we see that compared to baseline values, the number of filters in the first layer has little impact. The number of filters in layers 3-5 however are much more important no matter what resampling has been applied. The size of the filter in the first layer is interesting because it seems important with no resampling or little resampling, but becomes unimportant when resampling to 12 kHz or less.
4.2.4 Filtering
@@
When processing data sampled at 48 kHz we have nine filtering conditions: low-pass with a threshold of 12 kHz, 6 kHz, 3 kHz, 1.5 kHz, 750 Hz, 375 Hz, 187 Hz, 93 Hz and 46 Hz. These can be compared side by side. We use the methods described earlier, namely employing Shapley values and the d-dimensional earth mover’s scores, to investigate the impact of filtering. It had been suggested that a wide kernel is beneficial primarily because it filters out high-frequency noise. The filtering experiments make it possible to verify whether this really is the benefit granted by a wide kernel. Here, we include the results obtained when filtering CWRU data, applying different thresholds to decide how many frequencies to filter out.
Figure 11: Correlation between filtering conditions with 48 kHz CWRU data.
Correlations between the different conditions appear in Figure 11. We see that many levels of filtering, especially the weakest forms that only remove the highest frequencies, give correlations in the range 0.5 to 0.8. These quite high correlations imply that moderate filtering did not have a large impact on how to set the hyperparameters. Filtering at a threshold lower than 187 Hz however causes a sudden drop in correlations. The most extreme forms of filtering evidently do affect which hyperparameter settings work well. We observe that the average accuracy is much lower when filtering at 46 or 93 Hz (the average accuracy was 0.29 in both cases) than when filtering at 3000 Hz or above (the averages ranged from 0.95 to 0.97), and there might simply be no distinction between good and bad hyperparameter settings when using extreme filtering. If there are no good and bad hyperparameter settings then the results will be more influenced by noise and will correlate less, which is one possible explanation for the relatively low correlations seen. In summary, hyperparameters generalise fairly well and so the same settings do well on multiple levels of filtering, except when the strongest filtering is applied. Filtering changed the data properties but did not completely redefine which settings are good or bad.
@@
In order to make a deeper dive into the kernel size in the first layer we can plot the accuracy scores visually for each kernel size. Figure 12 shows the spread of accuracy scores obtained when using different lengths of kernel and when processing data with different levels of filtering. If the filtering is anything milder than 93 Hz we see that a larger kernel performs better. The value of a wide kernel does not relate to high-frequency noise as was suspected because any such noise would be removed by the filters. It seems that the benefit of a wide kernel must be explained in some other way.
Figure 13: Feature importance after filtering 48 kHz CWRU data at different cutoff thresholds. Each row is normalised to show the relative importance of the hyperparameters.
We derive Shapley values that represent how much each hyperparameter helps when predicting how accurate a network will be. These importance values are large when a hyperparameter really tells us something about how well a network will perform. Figure 13 shows Shapley values derived from networks trained to classify faults from vibration data with different levels of filtering. Both the size of the first filters and the number of filters in the later layers are important. Note that the balance of importance between these two hyperparameters changes as filtering is increasingly applied but it changes without any obvious pattern.
Figure 14: How much the accuracy scores change as a result of tweaking a hyperparameter, when processing 48 kHz CWRU data that was filtered at different cutoff thresholds. Each cell is normalised to show its value relative to surrogate data (see explanation in Section 3.3.6).
The earth mover’s distance (EMD) also paints a picture of how important different hyperparameters are. We show the EMD for our hyperparameters in Figure 14. This again confirms that the kernel size in the first layer and the number of filters in layers 3-5 are both important hyperparameters. The number of filters seems to become slightly more important when the filtering is increased.
5 Discussion: Hyperparameters of a Wide-Kernel CNN Explained
Having thoroughly analysed the hyperparameter space for our wide-kernel CNN we can now lay down some guidelines. We researched architectural hyperparameters such as the size of convolutional kernels in different layers and now we are ready to give the following advice. First, we note that some hyperparameters are less important and it should be possible to set these low in order to save computation. The kernel size and number of convolutional filters in the first layer are more important however and so we took additional care when analysing these. The number of filters in layers 3-5 is also important. We observe that larger kernels do better in the first layer and so suggest using a kernel of size 256 but we found that this depends on the dataset. Testing out different numbers of filters in layers 3-5 is a good way to improve performance although we see that a default value of 32 does well on multiple datasets.
Considering all of the above, we are then able to give fairly specific guidance on how to set the hyperparameters of the network. Going through the hyperparameters in turn:
  • 1.
    First layer kernel size: As a first attempt, it seems best to set this high and therefore use a wide kernel. Of all options we found that the highest value of 256 did the best usually. However, we also found that this hyperparameter is highly sensitive to differences between the datasets and so it should be carefully tuned if the initial accuracy of the network seems to leave room for improvement.
  • 2.
    Stride in the first layer: This does not greatly impact the accuracy so it may as well be set to the most computationally efficient value, which is actually the highest value since a long stride allows the network to skim faster over the inputs. In our experiments, that was 32.
  • 3.
    Number of filters in the first layer: We considered 8, 16, 32, 64, 128 and 256 as options and found that 16 and 32 are the best. We recommend 16 since it balances accuracy and computational efficiency. Although our results suggest it will not be necessary in most cases, since this is an important hyperparameter the user might decide to search for the optimal value amongst all the options.
  • 4.
    Second layer kernel size: This hyperparameter does not greatly impact the accuracy so it may as well be set to the most computationally efficient value. In our experiments, that was 3.
  • 5.
    Number of filters in the second layer: This hyperparameter also does not greatly impact the accuracy and can be set to the most computationally efficient value, which was 8 in our experiments.
  • 6.
    Kernel size in layers 3 to 5: This is another hyperparameter that does not greatly impact the accuracy and may as well be set to the most computationally efficient value. In our experiments, that was 3.
  • 7.
    Number of filters in the second layer: We considered 8, 16, 32, 64, 128 and 256 as options and found that 16 and 32 are the best. We recommend 32 as a reasonable default value. Although our results suggest that it will not be necessary in most cases, since this is an important hyperparameter the user might choose to search for the optimal value amongst all the options.
Our experiments suggest that if we use the default values recommended above, on average we would obtain an accuracy better than 75% of alternatives across the 7 datasets. Additionally tuning the kernel size in the first layer gives an accuracy that is better than 82% of alternative hyperparameter combinations, on average across all the datasets. It is worth noting also that increasing the number of filters in the first two layers to 32 gives a further boost so that the accuracy is higher than about 88% of alternative hyperparameter combinations. The number of filters in the first two layers seems to have increased importance in the specific context of a network tuned according to the advice we give above. In this context, the accuracy is better when using 32 and 32 instead of 16 and 8 filters in the first and second layers, providing another option for users to consider if they need to boost accuracy. Overall, putting our advice into practice appears to give competitive accuracy scores across multiple datasets.
6 Conclusion

### v2

@@ III-D Summary of Datasets Used
Across the various parts of our research, multiple datasets are employed. For the sake of clarity we now present the full collection of datasets we use and remark on which architectures they are used with. Table IV shows these details.
@@
TABLE IV: Summary of datasets used in the research.
IV Results
@@
IV-A Results Pt. 1 – Experiments on Multiple Architectures
Before we focus on a specific wide-kernel CNN architecture in part 2 of our results, we first present some experiments with two other popular architectural styles: a recurrent LSTM type of neural network and a transformer network. These results establish the general importance of hyperparameter tuning when working with neural networks.
IV-A1 Long Short Term Memory (LSTM)
Across our sample of 100 LSTM configurations, we see that poor hyperparameters can be disastrous. In the worst cases (CWRU and XJTU) the lowest-performing network was wrong more than 10 times as often as the best network. On all datasets the default values do roughly as well as about 75% of modified hyperparameter settings and therefore are reasonable choices. One interesting thing is that the LSTM network’s default values appear to under-perform on XJTU data. Facing this benchmark, an LSTM configured with the defaults only achieved an accuracy of 72% whilst it was possible to get above 90% with other hyperparameter values. Re-tuning the default hyperparameter settings from the original authors seems to be a good idea on XJTU data, indicating that hyperparameters that perform well on one dataset do not necessarily perform well on another dataset. The default values fared relatively much better on CWRU and SEU. Overall the results show that hyperparameter tuning has a noticeable impact on the LSTM architecture much like with the transformer and wide-kernel CNN networks.
Min 25% Median 75% Max Default
@@ IV-A1 Long Short Term Memory (LSTM)
TABLE V: Tukey’s five-number summary of how accurate the LSTM architecture was with different hyperparameter settings. The accuracy score when using the default hyperparameter values is also shown in the right-hand column.
IV-A2 Transformer
The transformer architecture performs relatively well when using the default hyperparameter settings, as shown in Table VI. On four of the baselines it was able to correctly identify fault conditions on more than 90% of test cases. Looking across all of the baseline datasets, the default hyperparameter settings were at least as good as 75% of altered versions of the settings. The results suggest that the default values are quite reasonable choices. On the Gearbox dataset the transformer was inaccurate no matter if the default hyperparameter settings were used or not, suggesting a specific problem with that one dataset. Importantly, our results show overall that the transformer architecture is sensitive to hyperparameter tweaking. On CWRU and UOC for example, the accuracy plunges by roughly 90% when swapping the best-performing hyperparameters for the worst.
@@ IV-A2 Transformer
TABLE VI: Tukey’s five-number summary of how accurate the transformer architecture was with different hyperparameter settings. The accuracy score when using the default hyperparameter values is also shown in the right-hand column.
IV-B Results Pt. 2 – Focusing on the Wide-Kernel CNN Architecture
The main body of our research is about the wide-kernel CNN architecture described in Section III-C1. First, we revisit how successful networks in a grid search are when trained and tested on seven different benchmark datasets. This grid search is the same as was analysed before in [1], although here our results have a different focus. We look at the average test results when setting individual hyperparameters to specific values, and we also zoom in on the interactions between the hyperparameters that were previously identified as most important. These results help us formulate guidance on how to tune the hyperparameters of the wide-kernel CNN. Following that, we also present multiple defaults, a small selection of the most effective hyperparameter combinations to try when moving to new data.
@@
IV-B1 Summarising Grid Search Results
Now we look at how the accuracy of the wide-kernel CNN varies as we change both the hyperparameter values and the data. Each condition’s accuracy is shown in Table VII. These results are based on the same grid search used previously in [1], which trained and tested different combinations of hyperparameters on 7 benchmarks. We revisit the grid search in order to pull out and highlight helpful information when thinking about how to tune the hyperparameters on different datasets.
@@ IV-B1 Summarising Grid Search Results
Kernel size 16 25 78 47 74 87 32 64
layer 1 32 31 78 48 84 80 35 59
@@
Kernel size 3 43 78 50 85 71 43 67
layer 2 6 45 77 49 84 70 43 62
Filters 8 44 78 52 85 74 49 68
layer 2 16 50 78 52 85 73 50 66
@@
Kernel size 3 45 80 50 84 71 43 65
layers 3-5 6 43 75 49 85 69 42 64
Filters 8 33 83 49 86 67 49 75
layers 3-5 16 45 91 50 89 72 51 82
@@
Mean accuracy 44 77 49 85 70 43 65
TABLE VII: The average performance for different values of wide-kernel CNN hyperparameters when trained and tested on different datasets.
@@
Perhaps more interesting are the individual accuracy scores showing what happens when we tune different hyperparameters. The kernel size in the first layer reveals the most variable results. With CWRU, Paderborn and UOC there is a strong effect whereby a longer kernel does better and the longest kernel exceeds the shortest by at least 15% accuracy. The MFPT and XJTU benchmarks exhibit a weaker effect of roughly 6% accuracy increase between the shortest and longest kernel. The kernel size in layer 1 barely effects the results on Gearbox data. Meanwhile, a strong contradictory effect is observed when using SEU data. The shortest kernel is much better (more than 30% better) compared to the longest kernel size. In all, it seems that longer kernels in the first layer tend to be best but it very much depends on the fault vibration data.
@@
Past work also identified the number of filters in the first layer as important. Good performance is often achieved when there are 16 or 32 filters. Either 16 or 32 was the best on every dataset except XJTU (when 8 filters was 1% better) and Gearbox. With Gearbox, the trend is that more filters gives higher accuracy and 256 beats 8 filters by about 7% accuracy. Although there are exceptions across the datasets, normally 16 or 32 are the safest bets when tuning this hyperparameter.
The final ‘important’ hyperparameter is the number of filters in layers 3 to 5. We observe that either 16 or 32 works best on all the benchmark datasets.
IV-B2 Influence of Important Hyperparameters on One Another
We now take another deep dive into the three hyperparameters identified as important by our previous work.
@@ IV-B2 Influence of Important Hyperparameters on One Another
It is very computationally expensive to run a full grid search over all the possible hyperparameter combinations and so someone applying the network to new data might well want to know how to tune the hyperparameters one-at-a-time. We ask which is the best order to tune the three most important hyperparameters sequentially. The method we used in our previous paper [1] comes in handy again here: we look at how much one hyperparameter influences another hyperparameter which has already been tuned. Because the hyperparameters are interconnected it is possible that a hyperparameter that started off being optimal no longer has the best value when a different hyperparameter changes. We look at how often this phenomenon happens and we summarise the information as an ‘influence’ score. If A is highly influential on B then changing A most likely means changing the best value for B. We can see the numbers in Figure 2.
@@
The kernel size in layer 1 appears to be the most influential and therefore it appears to be the best one to tune first. Tuning it later on would likely disrupt the other hyperparameters and cause them to need re-tuning. In other words, there is often no point in tuning the number of filters in layers 1 or 3-5 beforehand, since they will need to be tuned again when the kernel size changes. After choosing the size of filters in the first layer it is best to choose how many next. The choice of how many filters in layers 3 to 5 can be delayed to the end. In brief, the best order to tune one-at-a-time is: kernel size layer 1, number of filters layer 1, number of filters layers 3-5.
IV-B3 Multiple Defaults
Multiple defaults provide a sequence of hyperparameter combinations to try. The goal is that by trying them out sequentially, the user will be able to identify as quickly as possible at least one hyperparameter combination that works well on their data. This provides an alternative to tuning methods that are fast but do not consider interactions between hyperparameters (tuning one-at-a-time) or are slow but explore more options (e.g., grid search, random search, Bayesian optimisation). Multiple defaults are fast to use but also flexible and able to account for interactions between hyperparameters.
We used a brute force search to identify the global best series of hyperparameter combinations which terminated after finding 10 combinations. These are shown in Table VIII with the observed performance on the seven benchmarks.
No. KS 1 S 1 F 1 KS 2 F 2 KS 3+ F 3+ CWRU Gearbox MFPT Paderborn SEU UOC XJTU
1 256 16 256 3 16 3 64 0.96 0.89 0.93 0.88 0.57 0.99 0.96
2 16 16 128 6 128 3 64 0.08 1.00 0.07 0.26 0.97 0.12 0.58
3 256 8 32 3 32 6 64 0.98 0.68 0.99 0.98 0.16 0.98 0.99
4 16 4 8 6 32 3 8 0.44 0.40 0.61 0.35 1.00 0.46 0.69
5 128 16 32 6 64 6 128 0.97 0.39 1.00 1.00 0.35 0.84 0.49
6 256 4 64 6 128 6 128 1.00 0.00 0.22 0.29 0.07 0.40 0.15
7 256 4 256 6 8 3 16 0.74 0.94 0.66 0.95 0.51 0.60 1.00
8 256 16 32 6 256 6 128 0.51 0.16 1.00 0.91 0.13 1.00 0.55
9 256 16 64 6 64 3 64 0.96 0.85 0.97 0.54 0.25 1.00 0.43
10 32 8 32 3 8 6 128 0.51 0.23 1.00 0.69 0.80 0.22 0.25
TABLE VIII: How multiple defaults perform on the seven benchmark datasets.
We observe from Table VIII that the multiple defaults are complementary to each other, with the first entry performing relatively well on most benchmarks but having mediocre performance on the SEU dataset. The second entry picks up the slack and has very high performance on SEU. The third entry has excellent performance on the Paderborn dataset, improving on the previous two entries by 10%. The hyperparameter values of the first three entries tell a similar story. The first entry has the largest kernel size and many filters in the first layer. By contrast, the second entry has the smallest kernel size but many more filters in the second layer. The third entry also has the largest kernel size but combines it with a modest number of filters.
To illustrate how quickly these defaults identify a high-performing combination of hyperparameters, the average quantile (across benchmarks) of the best performer after trying $n$ hyperparameter combinations is visualised in Figure 3. We see that the expected performance increases quickly and there is already very little room to improve after trying three of the defaults. This gives hope that it will be possible to rapidly identify an effective combination of hyperparameters when working on new datasets. From a user’s perspective, a lot of time can be saved from the usually arduous process of hyperparameter tuning.
Figure 3: Expected performance relative to other hyperparameter settings, after trying multiple defaults. Calculated by taking the best-performing CNN tried so far per benchmark, then averaging the quantile of how well those CNNs did.
To get some idea of generalisation to new datasets, we repeatedly run and evaluate the search for multiple defaults using leave-one-out validation. The average best performance when generalising to an unseen benchmark is then computed. On average the multiple defaults get to the 92nd percentile on unseen data. Multiple defaults make it possible to identify hyperparameter combinations that perform well on new data, relative to the great majority of other hyperparameter combinations included in the full grid search.
IV-B4 Resampling
@@
The results of a grid search over the three most important hyperparameters, performed once each for the different resampling conditions, are brought together in the results here. Visualising the results makes it possible to spot any obvious trends in how the (increasing levels of) resampling impact the performance of different hyperparameter settings.
@@
Figure 4: Correlation between the same CNN configurations in terms of performance when using different resampling conditions, starting from 48 kHz.
@@
First, we consider how much the accuracy scores change as a consequence of resampling. Correlations tell us whether the best-performing combinations of hyperparameters on one version of the data also remain the best after resampling, and vice versa with the worst performers. The computed correlations are shown in Figure 4. All correlations are positive which implies that the hyperparameter values that do well on one version of the data have a tendency to perform well on data resampled to a different rate. The weakest correlation is between the original 48 kHz data and the most resampled data of 3 kHz. This is not surprising since they are the furthest away from each other in terms of data properties.
@@
Before moving on we must make a brief disclaimer about the correlations, specifically there is a ceiling effect because many networks achieve close to the maximum accuracy of 1.0, and, the training and testing of neural networks is stochastic and so there may be some random noise in the accuracy scores. There is a chance that these statistical quirks have inflated or deflated the correlations, and so we do not report p-values which might be biased by these effects.
@@
IV-B5 Filtering
@@
When processing data sampled at 48 kHz we have nine filtering conditions: low-pass with a threshold of 12 kHz, 6 kHz, 3 kHz, 1.5 kHz, 750 Hz, 375 Hz, 187 Hz, 93 Hz and 46 Hz. It had been suggested that a wide kernel is beneficial primarily because it filters out high-frequency noise. The filtering experiments make it possible to verify whether this really is the benefit granted by a wide kernel.
Figure 6: Correlation between the same CNN configurations in terms of performance when using different filtering conditions.
Correlations between the different conditions appear in Figure 6. We see that many levels of filtering, especially the weakest forms that only remove the highest frequencies, give correlations in the range 0.5 to 0.8. These quite high correlations imply that moderate filtering did not have a large impact on how to set the hyperparameters. Filtering at a threshold lower than 187 Hz however causes a sudden drop in correlations. The most extreme forms of filtering evidently do affect which hyperparameter settings work well. We observe that the average accuracy is much lower when filtering at 46 or 93 Hz (the average accuracy was 0.29 in both cases) than when filtering at 3000 Hz or above (the averages ranged from 0.95 to 0.97), and there might simply be no distinction between good and bad hyperparameter settings when using extreme filtering. If the meaningful content has been filtered out of the signal, only noise will remain and the performance of different hyperparameter combinations will be essentially random (thus uncorrelated). In summary, hyperparameters generalise fairly well and so the same settings do well on multiple levels of filtering, except when the strongest filtering is applied. Filtering changed the data properties but did not completely redefine which settings are good or bad.
@@
In order to make a deeper dive into the kernel size in the first layer we can plot the accuracy scores visually for each kernel size. Figure 7 shows the spread of accuracy scores obtained when using different lengths of kernel and when processing data with different levels of filtering. If the filtering is anything milder than 93 Hz we see that a larger kernel performs better. The value of a wide kernel does not relate to high-frequency noise as was suspected because any such noise would be removed by the filters. It seems that the benefit of a wide kernel must be explained in some other way. This is interesting because the ability to filter out high-frequency noise has been suggested as a reason why wide-kernel CNNs perform well.
V Conclusion